OpenStudy (anonymous):

what are the first five terms a1=0 and an=3.5-3a(n-1)

OpenStudy (accessdenied):

Have you tried anything so far? Or not sure how to begin?

OpenStudy (accessdenied):

The terms are defined in terms of the previous term in the sequence. We know the first term a1 = 0. If we want the second term, a2, try subbing n=2 into our formula: a2 = 3.5 - 3a(2-1) a2 = 3.5 - 3a1 We already defined a1, right? That was a1 = 0. So we can apply that to this formula. a2 = 3.5 - 3(0) = 3.5. What about a3? Apply the formula again, this time for n=3. We get something in terms of a2, which we just found. We plug a2 in, and then get a3. The process just repeats until we get all the values a1 through a5. Can you see how to proceed from here?